6. How many silver coins, 1.75 cm in diameter and of thickness 2 mm, must be melted to form a cuboid of dimensions 5.5 cm × 10 cm × 3.5 cm?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
It is known that the coins are cylindrical in shape. So, height (h1) of the cylinder = 2 mm = 0.2 cm Radius (r) of circular end of coins = 1.75/2 = 0.875 cm Now, the number of coins to be melted to form the required cuboids be “n” So, Volume of n coins = Volume of cuboids n × π × r2 × h1 = l × b × h n×π×(0.875)2×0.2 = 5.5×10×3.5 Or, n = 400